Consider the stock of Morgan Stanley., whose last dividend was $2.25. If the growth rate in dividends is 3.2% what is a fair price for Morgan Stanley stock if the market requires a 7.8% return on company's stock?
Answer #1

Ans $ 50.48

P0 = Price of Share
D1 = Current Dividend
Ke = Cost of Equity
g = growth rate
P0 = D1 / (Ke - g)
P0 = 2.3220 / (7.8%- 3.2%)
P0 = 50.48
D1 = D0* (1 + g)
D1 = 2.25* (1 + 3.2%)
D1 = 2.3220